OpenStudy (anonymous):

f(x) = - 2 cot 3x; what is the period and two vertical asymptotes? I believe the period is pi.

OpenStudy (anonymous):

\[\cot 3x= \frac{\cos 3x}{\sin 3x}\] asymptotes will occur when you divide by 0, so solve: \[\sin 3x = 0\]